Minutes of Working Group Meeting


IfA Hilo Facility, Hilo

June 12, 2014

Present:

Absent:

Minutes

  1. Current Network Status
  2. On April 28, starting around 10:00pm and lasting until noon on the 29th, issues with the UH backbone connections impacted all external connectivity out of Oahu. A temporary workaround was implemented which fixed IPv4 connectivity at ~10am on the 29th. Only commodity traffic was affected for the Big Island. Traffic through the SX South was not affected.

    According to Chris Zane of the UH-ITS, some routers were not receiving the multicast packet for OSPF hellos which caused the OSPF timer to expire, resulting in route flapping as OSPF adjacency went down. The problem eventually resolved itself. UH-ITS was not able to figure out the source of the problem.

    On May 20, for half an hour between 7:00am and 9:00am HST, due to scheduled maintenance that affected our two 10G circuits out of Oahu to the west coast as well as the SX North, all research traffic, other than that to and from the Big Island was rerouted through the Commodity Network. Since SX South was not affected, Big Island traffic was not affected. No issues were observed.

    On May 28, between 7:15am and 11:00am, there were some problems connecting to the Internet, which included access to Google@UH. These problems were resolved by around 11:00am, although the cause was not determined.

    On June 9, beginning early in the morning, there again were issues accessing the Internet, hawaii.edu websites, and UH email. The problems were traced to the Palo Alto device. All problems appeared to be fixed by ~11:45am.

    With two 10G circuits up and running for sometime between the HP and Hilo, UH-ITS is hoping to cancel the backup 1G HawTel circuit, for which they are paying. Pui Hin has asked to keep it for another six months -- as an alternate path for the observatories until the new Time Warner fiber path along the Saddle Road to West Hawaii is completed.

  3. FY 2014-15 Budget
  4. For this year, there are some changes to the budget as well as how the costs are distributed.

    In the past, all the observatories shared the annual cost equally. Starting this fiscal year, the directors have agreed to distribute the maintenance costs based on observatory size.

    Aside from the above, the total cost to be funded is much lower than last year for two reasons:

    1. The backup T1 circuit between HP and Manoa has been cancelled, bringing the Telephone/Telecommunications cost from $9,100 down to $2,500.

    2. Fund balance at the end of the 2013 fiscal year was -$9,057 due to the incident of the humidifier at the Summit being turned on and which consumed a lot of electricity. The projected fund balance this year is $8,227.

    The above two factors are responsible for bringing the total maintenance cost to be funded from $109,661 last year down to $77,477 this year. All other changes are small adjustments of no consequence.

  5. Upgrade of the Observatories' Connections to the MKOCN
  6. Pui Hin reported that SMA has migrated to 10Gb on the Summit and in Hilo. The migration was fairly easy and without incident, although the alternate routes still need some work with UH-ITS to get it going. It is worth noting that in Hilo, Mac is able to get 10G working over the OM2 500MHz 50/125 multimode fiber installed in 2009. According to Cisco datasheets, SFP-10g-LRM will support up to 220m at 1310nm over 500MHz OM2 fiber cable. You can get the AddON brand equivalence for ~$350 each.

  7. Weather Display Station at HP
  8. All equipment has been ordered.

  9. Next meeting
  10. March 31st, Tuesday, 10:00am at the IfA Manoa Facility.


MKOCN Homepage Back to Minutes

Last updated March 30, 2015 by Miranda Hawarden-Ogata